OpenStudy (gabylovesyou):

During a science experiment, Faye recorded the temperature of a solution as it cooled in a controlled environment for 6 minutes. She graphed the change in temperature as shown below.

OpenStudy (gabylovesyou):

If the solution continues to cool at the same rate, how many minutes will it take for the temperature to cool to 0°C? 24 19 14 9

OpenStudy (gabylovesyou):

whoops wrong pic

OpenStudy (gabylovesyou):

@mathteacher1729 @math_proof @mathteacher1729

OpenStudy (mathteacher1729):

How do they want you to solve this? (graphing, algebra, estimation?) Can you use a calculator?

OpenStudy (gabylovesyou):

yea a calculator

OpenStudy (gabylovesyou):

its A

OpenStudy (gabylovesyou):

right

OpenStudy (mathteacher1729):

Sorry I was drawing a graph...

OpenStudy (mathteacher1729):

Oh sorry, it should say right three MINUTES not days. And yes if you solve 0 = (-5/3)x + 40 add (5/3)x to both sides (5/3)x = 40 multiply both sides by (3/5) x = 40*(3/5) x = 24 minutes So you're correct. :)